# Do LLMs “know” internally when they follow instructions?

### v4

@@ Abstract
Instruction-following is crucial for building AI agents with large language models (LLMs), as these models must adhere strictly to user-provided constraints and guidelines. However, LLMs often fail to follow even simple and clear instructions. To improve instruction-following behavior and prevent undesirable outputs, a deeper understanding of how LLMs’ internal states relate to these outcomes is required. Our analysis of LLM internal states reveal a dimension in the input embedding space linked to successful instruction-following. We demonstrate that modifying representations along this dimension improves instruction-following success rates compared to random changes, without compromising response quality. Further investigation reveals that this dimension is more closely related to the phrasing of prompts rather than the inherent difficulty of the task or instructions. This discovery also suggests explanations for why LLMs sometimes fail to follow clear instructions and why prompt engineering is often effective, even when the content remains largely unchanged. This work provides insight into the internal workings of LLMs’ instruction-following, paving the way for reliable LLM agents.111We will release the data and code on GitHub upon publication.
@@ 1 Introduction
However, LLMs often fail to follow even non-ambiguous and simple instructions (Zhou et al., 2023; Qin et al., 2024; Xia et al., 2024; Kim et al., 2024; Yan et al., 2024) like including keywords or following formatting guidelines. GPT-4 achieves around an 80% success rate on IFEval (Zhou et al., 2023), an instruction-following benchmark dataset, while smaller models have success rates around 30% to 40%. This raises the question: why do LLMs fail to follow instructions, even when those instructions are clear and familiar?
To gain a better understanding of instruction-following outcomes, we analyze the internal state of LLMs, focusing on the differences in representations between success and failure cases of instruction-following across different tokens and layers. Our approach involves disentangling the effects of tasks and instructions in input prompts, where the instruction specifies the action (e.g., ‘please do not use keywords’) and the task provides the context for executing the instruction (e.g., ‘please write a resume’). By applying linear probing—a widely used method for interpreting model representations (Alain & Bengio, 2016; Belinkov, 2022; Elazar et al., 2021)—we identify a specific dimension within the input embedding space that is strongly associated with instruction-following. While previous work has primarily used linear probing to explore representations related to truthfulness and reducing hallucinations (Azaria & Mitchell, 2023; Marks & Tegmark, 2023; MacDiarmid et al., 2024), our study extends this method to investigate instruction-following. We demonstrate that this dimension generalizes to unseen tasks, indicating that it captures a fundamental aspect of instruction adherence in LLMs.
@@ 2.1 IFEval-simple
First, we selected the IFEval dataset (Zhou et al., 2023) as a base, due to its objective evaluations with verifiable instructions, thereby minimizing uncertainties from ambiguous evaluation criteria. The IFEval dataset comprises 25 instruction types under 9 categories, with each instruction type paired with a distinct set of tasks — approximately 20 tasks per instruction type. Because of the relatively small number of tasks per instruction type, internal model states resulting from these prompts contain a mix of both instruction-following and task-specific details.
To isolate the dimension related specifically to instruction-following, we generated a modified version of the IFEval data, called IFEval-simple.222We will release the IFEval-simple data along with code on GitHub upon publication. First, we selected 5 instruction types that are likely to be used in real-world applications for AI agents. For example, ensuring the inclusion (keywords:existence) or exclusion (keywords:forbidden) of specific keywords, specifying the frequency of certain keywords (keywords:frequency), generating responses with placeholders (detectable_content:place_holders), and requiring responses to end with predefined sentences (startend:end checker). We excluded more complex or impractical instructions, such as those requiring omission of punctuation, as they are less relevant for practical use cases.
@@ 2.2 Methods
Representations We analyzed four language models: LLaMA-2-7B-chat (Touvron et al., 2023), LLaMA-2-13B-chat (Touvron et al., 2023), Mistral-7B-Instruct-v0.3 (Jiang et al., 2023), and Phi-3-mini-128k-instruct (Abdin et al., 2024). For each model, we looked at the representations between tokens – the first, middle, and last tokens, representing the LLMs before, during, and after they generate responses. We also examined three layers (early, middle, last) to identify where instruction-following information is encoded within the models’ internal state. Specifically, we used layers 16, 32, and 40 and for LLaMA-2-13B-chat and 14, 26, and 32 for other three models. To avoid randomness in decoding, we employed greedy decoding without sampling.
@@ 4 Interpreting the instruction-following dimension
(2) Instruction Difficulty: We investigated the relationship of the instruction-following dimension with the complexity of the instructions. We perturbed the instruction difficulty by simplifying instructions by relaxing instruction-related constraints. For example, in the original instruction “please include keywords: coding, Python, computer, experience”, we reduced the complexity by reducing the number of keywords required in the instruction to “please include the keywords: coding, Python”.
@@ 5 Related work
Instruction-following in LLMs Recent research has introduced various benchmark datasets to evaluate the instruction-following capabilities of LLMs across different contexts(Zhou et al., 2023; Qin et al., 2024; Yan et al., 2024; Xia et al., 2024). Among these, we selected the IFEval dataset (Zhou et al., 2023) as the foundation for our study, due to its structured and objective evaluation framework that minimizes ambiguity by using verifiable instructions. Beyond evaluation, several approaches have been proposed to improve instruction-following performance, such as modifying attention mechanisms (Zhang et al., 2023) and applying fine-tuning strategies (He et al., 2024; Sun et al., 2024). In contrast to prior work that primarily focuses on evaluating or enhancing instruction-following, our study aims to understand why LLMs sometimes fail to follow instructions by analyzing their internal representations.
@@ 6.1 LLMs internally recognize whether they will follow instructions
LLMs generalize well across tasks but struggle with different instruction types We find that while LLMs can generalize across different tasks, they struggle with generalization across different instruction types. This suggests that distinct instruction categories may have unique geometries within the models’ internal representation space, making it more challenging for the model to generalize across them.
@@ A.4 Success rate
A.5 More result on Representation Engineering
In the main paper, we presented the average performance of Representation Engineering (RE) across all instruction types and models. Here, we provide more detailed insights through figures, highlighting the variation in RE effectiveness across different types of instructions.
Figure 3 shows the transition metrics for high-quality responses, including F2T (False to True) and T2T (True to True). These results demonstrate the robustness of the instruction-following dimension in improving success rates. The proportion of F2T and T2T transitions is consistently higher when using the instruction-following dimension compared to random directions across various instruction types. This indicates that RE not only enhances success rates but also effectively converts failed responses into successful ones, while maintaining overall task quality.
A.6 PCA across all five instruction types

### v5

@@ Abstract
Instruction-following is crucial for building AI agents with large language models (LLMs), as these models must adhere strictly to user-provided constraints and guidelines. However, LLMs often fail to follow even simple and clear instructions. To improve instruction-following behavior and prevent undesirable outputs, a deeper understanding of how LLMs’ internal states relate to these outcomes is required. In this work, we investigate whether LLMs encode information in their representations that correlates with instruction-following success—a property we term “knowing internally”. Our analysis identifies a direction in the input embedding space, termed the instruction-following dimension, that predicts whether a response will comply with a given instruction. We find that this dimension generalizes well across unseen tasks but not across unseen instruction types. We demonstrate that modifying representations along this dimension improves instruction-following success rates compared to random changes, without compromising response quality. Further investigation reveals that this dimension is more closely related to the phrasing of prompts rather than the inherent difficulty of the task or instructions. This work provides insight into the internal workings of LLMs’ instruction-following, paving the way for reliable LLM agents.111Code and data are available at https://github.com/apple/ml-internal-llms-instruction-following
@@ 1 Introduction
However, LLMs often fail to follow even unambiguous and simple instructions (Zhou et al., 2023; Qin et al., 2024; Xia et al., 2024; Kim et al., 2024; Yan et al., 2024) like including keywords or following formatting guidelines. GPT-4 achieves around an 80% success rate on IFEval (Zhou et al., 2023), an instruction-following benchmark dataset, while smaller models have success rates around 30% to 40%. This raises the question: why do LLMs fail to follow instructions, even when those instructions are clear and familiar?
To gain a better understanding of instruction-following outcomes, we analyze the internal state of LLMs, focusing on the differences in representations between success and failure cases of instruction-following across different tokens and layers. Our approach involves disentangling the effects of tasks and instructions in input prompts, where the instruction specifies the action (e.g., ‘please do not use keywords’) and the task provides the context for executing the instruction (e.g., ‘please write a resume’). By applying linear probing—a widely used method for interpreting model representations (Alain & Bengio, 2016; Belinkov, 2022; Elazar et al., 2021)—we identify a specific dimension within the input embedding space that is strongly associated with instruction-following. While previous work has primarily used linear probing to explore representations related to truthfulness and reducing hallucinations (Azaria & Mitchell, 2023; Marks & Tegmark, 2023; MacDiarmid et al., 2024), our study extends this method to investigate instruction-following. We demonstrate that this dimension generalizes to unseen tasks, however not to unseen instruction types.
@@ 2.1 IFEval-simple
To objectively evaluate LLMs with simple and verifiable instructions, we select IFEval (Zhou et al., 2023) as our base dataset. The motivation is that, while complex and multi-purpose instruction prompts are more realistic, they require using LLM-based evaluators that may induce further errors and biases in assessing success or failure. To avoid this potential issue, we focus on simple, single-purpose and verifiable instructions from IFEval, such as “Please do not include keywords: …” or “answer in lower-case only”, that can be automatically validated with deterministic programs like string-matching, thereby minimizing uncertainties from ambiguous evaluation criteria. We provide a more detailed justification in Appendix A.6.
The IFEval dataset comprises 25 instruction types under 9 categories, with each instruction type paired with a distinct set of tasks — approximately 20 tasks per instruction type. Furthermore, due to the relatively small number of tasks per instruction type, internal model states resulting from these prompts contain a mix of both instruction-following and task-specific details. To isolate the dimension related specifically to instruction-following, we generated a modified version of the IFEval data, called IFEval-simple.222The IFEval-simple data is available at https://github.com/apple/ml-internal-llms-instruction-following. First, we selected 5 instruction types that are likely to be used in real-world applications for AI agents. For example, ensuring the inclusion (keywords:existence) or exclusion (keywords:forbidden) of specific keywords, specifying the frequency of certain keywords (keywords:frequency), generating responses with placeholders (detectable_content:place_holders), and requiring responses to end with predefined sentences (startend:end checker). We excluded more complex or impractical instructions, such as those requiring omission of punctuation, as they are less relevant for practical use cases.
@@ 2.2 Methods
Representations We analyzed four language models: LLaMA-2-7B-chat (Touvron et al., 2023), LLaMA-2-13B-chat (Touvron et al., 2023), Mistral-7B-Instruct-v0.3 (Jiang et al., 2023), and Phi-3-mini-128k-instruct (Abdin et al., 2024). For each model, we looked at the representations on three tokens: (1) first token, , where are the tokens in the input prompt; (2) middle token, , where are the first tokens of the response; and (3) last token, , representing the full input and response. We also examined three layers (early, middle, last) to identify where instruction-following information is encoded within the models’ internal state. Specifically, we used layers 16, 32, and 40 and for LLaMA-2-13B-chat and 14, 26, and 32 for other three models. To avoid randomness in decoding, we employed greedy decoding without sampling.
@@ 4 Interpreting the instruction-following dimension
(2) Instruction Difficulty: We investigated the relationship of the instruction-following dimension with the complexity of the instructions. We perturbed the instruction difficulty by simplifying instructions by relaxing instruction-related constraints. For example, in the original instruction “Please include keywords: coding, Python, computer, experience”, we reduced the complexity by reducing the number of keywords required in the instruction to “Please include the keywords: coding”.
@@ 5 Related work
Instruction-following in LLMs Recent research has introduced various benchmark datasets to evaluate the instruction-following capabilities of LLMs across different contexts(Zhou et al., 2023; Qin et al., 2024; Yan et al., 2024; Xia et al., 2024). Beyond evaluation, several approaches have been proposed to improve instruction-following performance, such as modifying attention mechanisms (Zhang et al., 2023) and applying fine-tuning strategies (He et al., 2024; Sun et al., 2024). In contrast to prior work that primarily focuses on evaluating or enhancing instruction-following, our study aims to understand why LLMs sometimes fail to follow instructions by analyzing internal representations.
@@ 6.1 LLMs internally recognize whether they will follow instructions
LLMs generalize well across tasks but struggle with different instruction types We find that while LLMs can generalize across different tasks, they struggle with generalization across different instruction types. This suggests that distinct instruction categories may have unique geometries within the models’ internal representation space, making it more challenging to generalize across them.
@@ 6.3 Limitations and Future work
Acknowledgments
This work was conducted during an internship at Apple AIML. We sincerely thank Fahad Kamran and Feng Zhu for their valuable feedback and insightful suggestions on this work. We are also grateful to Guillermo Sapiro for his unwavering support and guidance throughout the research.
@@ A.4 Success rate
A.5 PCA across all five instruction types
@@
A.6 Why do we choose IFEval dataset?
Here, we would like to emphasize why we choose IFEval as our primary dataset instead of using real-world dataset with different contexts and domains.
First, we select IFEval to focus on our scope which is ‘single, simple, and non-ambiguous instructions’. Real-world datasets often involve complex, ambiguous, or multi-instruction prompts, which can conflate multiple factors affecting instruction-following. As an initial exploration of the geometry of LLM representations in instruction-following, we chose to focus on single, simple, and verifiable instructions to ensure clarity and disentangle multiple factors. The IFEval dataset is well-suited for this purpose, as it provides 25 distinct types of simple and clear instructions that align with our goal of establishing a robust baseline.
Second, we want to avoid evaluator-induced uncertainties. Most real-world tasks and benchmark datasets rely on LLM-based evaluators to determine whether a response follows an instruction. However, LLM-based evaluators may introduce their own uncertainties or make errors in assessing success or failure, which could obscure our analysis on representations of the tested models. The IFEval dataset avoids this issue by including instructions with deterministic evaluation programs that objectively verify compliance. For instance, an instruction like “please do not include keywords: …” can be automatically validated using a simple program to check for the presence of those keywords. This feature eliminates ambiguity in evaluation and allows us to isolate the directions related specifically to instruction-following.
One of our main contribution is the careful design of data settings specifically tailored to analyze internal states of LLMs in instruction-following contexts. While IFEval serves as an ideal starting point for this research, we hope our work inspires future efforts to tackle analysis of LLMs in more complex, real-world instruction-following tasks.
A.7 Reverse Representation Engineering
We conducted initial experiments on reverse representation engineering with two models: Phi-3-mini-128k and Mistral-7B-inst-v0.3. In these tests, we try to move representations towards the failure class by flipping the adjustment vector
| Model | Original SR | Random SR | Reverse Inst-follow SR |
| Mistral | 0.58 ± 0.00 | 0.56 ± 0.02 | 0.54 ± 0.01 |
| Phi | 0.71 ± 0.00 | 0.63 ± 0.04 | 0.60 ± 0.02 |
Notably, we set the values conservatively to keep the quality ratio (QR) of reverse RE remains similar to that of random directions (0.86 for Mistral and 0.77 for Phi). The results indicate that the success rate (SR) for reverse RE is worse than random directions, as expected, but the difference is not significant. We anticipate that finding on a validation set will amplify the difference between reverse and random directions. We plan to conduct additional experiments to refine and better evaluate the effectiveness of reverse RE in disrupting instruction adherence.